A race car driver can complete 1 lap across a 2-mile track in 1 minute. Due to a slight malfunction, the car travels 40% slower. How long would it take the car to complete 1 lap with the malfunction?

Answer by stanbon(75887)

A race car driver can complete 1 lap across a 2-mile track in 1 minute. Due to a slight malfunction, the car travels 40% slower. How long would it take the car to complete 1 lap with the malfunction?
------
Before malfunction DATA:
distance = 2 miles ; time = 1 min ; rate = 2 mi/(1/60) hr = 120 mph
------------------
After malfunction DATA:
distance = 2 miles ; rate = 0.6*120 = 72 mph
time = d/r = 2 miles/ 72 mph = (1/36) hr = (1/36)(60) = 1 2/3 minutes
--------------------
Cheers,
Stan H.